The chances a person hits a target are 3/4 while the chances of another person hitting the same target are 2/3. Find the probability of the target being hit when they both try. Please provide detailed explanation.
Answer by stanbon(75887)

The chances a person hits a target are 3/4 while the chances of another person hitting the same target are 2/3. Find the probability of the target being hit when they both try. Please provide detailed explanation.
------
The 2 events are independent.
P(both hit target) = P(A hits)*P(Bhits) = (3/4)(2/3) = 2/4 = 1/2
------------
Cheers,
Stan H.